# UniT: Unified Tactile Representation for Robot Learning

### v1

@@ IV Tactile Perception Experiments
We use the following loss function to train the model, which converts quaternions to angle differences $L=\mathbb{E}\left[2\cdot\text{arccos}\left(\left|\hat{q}\cdot q\right|\right)\right],$ where $\hat{q}$ and $q$ are estimated and groundtruth quaternions. For more details on this experiment, please refer to our open-source repository [^3] https://github.com/ZhengtongXu/UniT.
The benchmark results are shown in Tables I and II. For representation learning methods, we freeze the encoders of UniT, MAE, BYOL, and T3 when implementing them to this 3D perception task. This approach allows us to evaluate whether the representation learning methods can truly learn generalizable and transferable tactile representations. The capability for zero-shot transfer is important when applying tactile representations across various perception and policy learning tasks with different objects (see Section V), as fine-tuning for each task and each object can be prohibitively costly.
Our experimental results clearly show that UniT delivered the best performance on this 3D pose estimation task. Moreover, compared to training from scratch, representation learning methods generally improve the performance of tactile perception. Finally, the experimental results also demonstrate that VQ regularization enhances the performance of downstream tasks. This indicates that a more compact tactile representation is beneficial for learning downstream tasks.
V Policy Learning Experiments
In this section, the effectiveness of UniT in imitation learning is demonstrated. We use Aloha [3] as our hardware platform. The hardware setup is shown in Fig. 6(a). We integrate UniT into diffusion policy [2], and benchmark with vision-only diffusion policy and visual-tactile diffusion policy with tactile encoders trained from scratch. For training datails, please refer to our open-source repository. We completed three tasks, as shown in Fig. 6. More details of each task are as follows.
@@ V Policy Learning Experiments
Chicken Legs Hanging: The chicken legs hanging task is a dual-arm manipulation task in which one robotic arm grasps the rack to prevent it from moving due to the pushing force exerted while hanging chicken legs. The other arm picks up two artificial chicken legs from a table and hangs them on the rack. The precision required for this task is very high, because the width of the chicken legs and the opening of the slots on the rack are almost a perfect match. This precise alignment ensures that the chicken legs remain securely in place on the rack after being released. Moreover, this task involves rich force interactions with the objects being manipulated: the left arm applies force on the rack and stabilizes it, while the right arm exerts pressure to insert the chicken leg into the slot on the rack. In this situation, real-time tactile feedback is crucial for fully perceiving these interactions.
@@
Allen Key Insertion: The Allen key insertion task requires the robot to grasp a Allen Key on a rack and then insert the Allen Key into a nut. We used a 6 mm nut and an Allen key with 5.5 mm head. In this configuration, the Allen key can still be used to tighten and loosen the nut effectively after insertion. Therefore, the precision requirement for this insertion is at the millimeter level. Given that Aloha is a low-cost robot, its repeatability in terms of precision is not high, which adds significant challenges to this task.
@@
During data collection and policy rollout, we modify the size of the rack (the height difference between the supports on each side), which in turn affects the angle of the Allen key when it is grasped in-hand by the robot. We collected data for two different types of racks in the training set. During policy rollout, we tested three different types of rack, including a rack that was completely unseen in the training set, representing out-of-distribution data (see Fig. 6(c)).
The tactile modality serves two pivotal functions in this task: 1) Given the variety of racks, leveraging tactile modality to gather information on the Allen key’s in-hand pose is vital for appropriately adjusting the pose of the end effector during insertion. 2) Insertion often proves challenging when aligning the Allen key with the nut on the first attempt, often leading to minor misalignments. In these instances, tactile feedback plays a critical role in interactively refining the policy to accurately guide the Allen key into the nut.
@@
The results of the policy rollout on these three tasks are shown in Table III. It is evident that the visual-tactile policy with UniT delivers the best performance on each task. As described earlier, the three tasks presented here involve rich robot-object-environment interactions. Our experimental results highlight two important points regarding these types of tasks: 1) The tactile modality is crucial for manipulation tasks with rich interactions, where relying solely on vision often fails to deliver optimal performance. 2) Compared to treating tactile images as regular visual images and training encoders from scratch, using UniT significantly enhances the effectiveness of integrating the tactile modality into the policy.
For more detailed information about our experimental results, see the supplementary video and the project website [^4] https://zhengtongxu.github.io/unifiedtactile.github.io/.
VI Discussion and Future Work
Due to the unique nature of tactile images, traditional visual representation learning methods may not be entirely suitable for tactile images. For example, in MAE, much of the masked input is ineffective tactile background because the useful information in a tactile image is primarily located where the object contacts the sensor pad. Inspired by the compact color distribution of tactile images, we utilized VQVAE to learn a compact latent space. This approach has been shown to be effective for representation learning specifically tailored to tactile images. Finally, two promising research directions are proposed:
1. Extension to Soft Objects: Our current framework has been validated only on rigid objects. However, the tactile features of soft objects pose a new set of challenges due to their complexity [11]. Exploring how to learn a unified tactile representation that works effectively for both rigid and soft objects should be a promising topic.
2. Physics-Informed Tactile Representations: Soft objects often correlate with complex physical properties, such as stiffness [11]. In addition, other studies have highlighted the potential of tactile sensing in understanding the physical world [12, 13]. Researching how to develop physics-informed representations that can integrate with diverse tasks (such as estimation of physical properties, physics-inspired robot manipulation, and AI4Science) could be highly intriguing.
VII acknowledgements
This work was partially supported by the United States Department of Agriculture (USDA; No. 2023-67021-39072 and No. 2024-67021-42878). This article solely reflects the opinions and conclusions of its authors and not of USDA.

### v2

@@ VI-A Pose Estimation
We use the following loss function to train the model, which converts quaternions to angle differences $L=\mathbb{E}\left[2\cdot\text{arccos}\left(\left|\hat{q}\cdot q\right|\right)\right],$ where $\hat{q}$ and $q$ are estimated and z quaternions. The loss function for 6D pose estimation includes quaternion-based angular error loss, supplemented with an additional MSE error loss for linear position. For more details, please refer to our open-source repository https://github.com/ZhengtongXu/UniT.
The benchmark results of 3D and 6D pose estimation tasks are shown in Tables I and II. A. K. represents Allen key dataset and S. B. represents small ball dataset. For 6D pose estimation, we test representations trained with Allen key data. For T3, we utilize the pretrained model released by the authors of [10]. We freeze the encoders of BYOL, T3, MAE, and UniT and train only the decoders on the pose estimation datasets. In the ResNet experiments, we train both the encoder and the decoders on pose data directly, as both ResNet and ResNet with pretrain do not go through a tactile representation learning process. The MAE with ViT-Base backbone has a model size of 143.20 M, whereas UniT with a representation dimension of $16\times 20$ has a model size of 79.81 M. Our experimental results clearly show that UniT delivered the best performance on both 3D pose and 6D pose estimation tasks.
Finally, the ablation study for UniT in Table I demonstrates the contributions of VQ and discriminator to learning effective tactile representations. Removing VQ results in worse performance across all tested representation dimensions and for both the Allen key and small ball representations. For discriminator, except for the small ball with an $8\times 10$ representation, most cases show better results when discriminator is included. Even in the small ball with $8\times 10$ representation case, the performance with and without discriminator is quite similar. Overall, including discriminator generally enhances model performance.
VI-B Classification
Our classification experiment uses the real-environment tactile data from the YCB-Sight dataset [29]. The authors of YCB-Sight have released a visuo-tactile dataset including both simulation and real data. The real environment data includes data from six objects, each with 40 tactile images with contact with the object. We used the 240 contact images to build our classification dataset, labeling each image by the object it contacted. We shuffled these 240 images and split them into training and test sets in an 8:2 ratio. The benchmarking results are shown in Table III. We report classification accuracy on the test set as our evaluation metric. “Freeze” means freezing the encoder and using the representation in a zero-shot manner to train the decoder on the classification training set. “Fine-Tune” means training both the decoder and fine-tuning the encoder on the classification training set. We tested two setups: one where the representation was transferred to this task in a zero-shot manner, and another where the encoders were fine-tuned specifically for this task. Our experimental results clearly show that UniT delivered the best performance in the zero-shot setup. When fine-tuned, UniT’s performance was comparable to that of ResNet with pretraining and BYOL. The strong performance of ResNet with pretraining and BYOL in the fine-tuning setup can be attributed to the small size of the dataset, as BYOL’s backbone is also based on ResNet. On the other hand, for pose estimation, ResNet’s performance is not as strong as UniT’s, as shown in Tables I and II.
Moreover, Table IV shows that, for UniT, a well-performing representation can be achieved with only a small amount of single-object data, allowing the trained representation to generalize in a zero-shot manner to unseen objects. This is similar to our use of the Allen key and small ball in pose estimation. By comparing Tables III and IV, UniT representations trained on one object like the sugar box or bleach cleanser achieve better zero-shot classification results than MAE and T3, even after fine-tuning. Notably, MAE representations here are trained on all six objects and T3 are pretrained on a large amount of tactile images.
VII Supervised Policy Learning Experiments
In this section, the effectiveness of UniT in imitation learning is demonstrated. We use Aloha [2] as our hardware platform. The hardware setup is shown in Fig. 7(a). We integrate UniT into diffusion policy [1], and benchmark with vision-only diffusion policy and visual-tactile diffusion policy with tactile encoders trained from scratch. We freeze the UniT encoder when integrating it into diffusion policy, as shown in Fig. 2. All three methods are trained on the same dataset, with each method using a diffusion policy as the main network, differing only in the structure of the tactile encoder, while the rest of the network architecture and hyperparameters are identical. We completed three tasks, as shown in Fig. 7. The “Chicken Legs Hanging”, “Chips Grasping”, and “Allen Key Insertion” datasets contain 200, 180, and 170 demonstrations, respectively. The action space is defined in terms of joint angles: each arm of the ALOHA dual-arm robot has 6 degrees of freedom, plus an additional degree of freedom for grasping width, totaling 7 degrees of freedom per arm. For the “Allen Key Insertion” task, which involves only one arm, the policy outputs 7-DoF actions. In contrast, both the “Chips Grasping” and ‘Chicken Legs Hanging” tasks are bimanual manipulation, so the policy outputs 14-DoF actions. The frequency of all policy rollouts is 10 Hz and the amount of inference denoising steps is 24. More details of each task are as follows.
@@ VII Supervised Policy Learning Experiments
Chicken Legs Hanging: The chicken legs hanging task is a dual-arm manipulation task in which one robotic arm grasps the rack to prevent it from moving due to the pushing force exerted while hanging chicken legs. The other arm picks up two artificial chicken legs from a table and hangs them on the rack. The rack slot has a trapezoidal shape, with a width ranging from 15.6 mm on the innermost side to 20.8 mm on the outermost side. In comparison, the chicken drumstick root varies in width from 19.2 mm at its narrowest to 22.2 mm at its widest. This close dimensional match necessitates precise alignment to securely hang the drumstick, underscoring the high level of precision required for successful task execution. Moreover, this task involves rich force interactions with the objects being manipulated: the left arm applies force on the rack and stabilizes it, while the right arm exerts pressure to insert the chicken leg into the slot on the rack. In this situation, real-time tactile feedback is crucial for fully perceiving these interactions.
@@
Allen Key Insertion: The Allen key insertion task requires the robot to grasp a Allen key on a rack and then insert the Allen key into a nut. We used a 6 mm nut and an Allen key with 5.5 mm head. Therefore, the precision requirement for this insertion is at the millimeter level. In this configuration, the Allen key can still be used to tighten and loosen the nut effectively after insertion. Given that Aloha is a low-cost robot, its repeatability in terms of precision is not high, which adds significant challenges to this task.
@@
During data collection and policy rollout, we adjusted the height difference between the rack’s supports, altering the Allen key’s in-hand angle when grasped by the robot. Data was collected for two rack types in the training set, while policy rollout tested three types, including one entirely unseen rack representing out-of-distribution data, as shown in Fig. 7(c).
The tactile modality plays a crucial role in this task: (1) It provides information on the Allen key’s in-hand pose, enabling precise end-effector adjustments across varying rack types during insertion. (2) When initial alignments fail, tactile feedback is essential for interactively refining the policy to guide the Allen key into the nut accurately.
@@
The results of the policy rollout on these three tasks are shown in Table V. It is evident that the visual-tactile policy with UniT delivers the best performance on each task. As described earlier, the three tasks presented here involve rich robot-object-environment interactions. Our experimental results highlight two important points regarding these types of tasks: 1) The tactile modality is crucial for manipulation tasks with rich interactions, where relying solely on vision often fails to deliver optimal performance. 2) Compared to treating tactile images as regular visual images and training encoders from scratch, using UniT significantly enhances the effectiveness of integrating the tactile modality into the policy. For more detailed information about our experimental results, see the supplementary video and the project website https://zhengtongxu.github.io/unit-website/.
Besides experiments in real environment, we constructed a simulated peg insertion task using TacSL [30], as shown in Fig. 8. To collect training data, we gathered 30 episodes through human demonstration with SpaceMouse. As shown in Fig. 8, the observation provides two camera images from different viewpoints and one tactile image from GelSight. The action space consists of six-degree-of-freedom end-effector velocity control. During both data collection and policy rollout, the robot’s end effector is initialized in different configurations, and the hole position is randomized within a -5 cm to 5 cm range along the x and y-axes. During rollout, task success is determined by checking the spatial relationship between the peg and the socket.
We train all policies using the same learning rate scheduler for 500 epochs. Every 25 epochs, we perform rollouts across 50 different environment initializations to measure the success rate. The final metric is computed as the average success rate over the rollouts from the last 10 checkpoints. Moreover, we repeat the evaluation across three different seeds and report the average results in Table VI. Our results demonstrate that policies with UniT outperform those with T3, as well as vision-only policies and visual-tactile policies trained from scratch.
VIII Future Work
For future work, we believe data-efficient tactile representation learning can be further improved by incorporating a diverse set of objects with fine textures to enhance generalization and performance. Specifically, investigating how training with a broader range of physical objects, including those with intricate fabric-like patterns, can enrich the representation’s ability to capture subtle tactile details is a promising direction.